Solve for x:
4x + 2y =14
x=3y
Please show steps. Thank you.

Answer by stanbon(75887)

Solve for x:
4x + 2y =14
x=3y
----
Modify:
2x + y = 7
y = (1/3)x
----
Substitute:
2x + (1/3)x = 7
(7/3)x = 7
x = 3
-------------------
Cheers,
Stan H.